OpenStudy (anonymous):

Which of the following describes the set {1,2,3, …}?

Parth (parthkohli):

There are a lot of notations for sets, one of them being \(\{ n\}_{1 \le n \le +\infty}\)

Parth (parthkohli):

If you want a description, this is the set of natural numbers.